# Growing Mathlib: maintenance of a large scale mathematical library

### v1

@@ 7 Scaling code review
Since its beginning, the Mathlib library required code review prior to merging changes. This is enforced through the bors bot: pull requests are merged automatically upon approval by an authorised user (i.e., a Mathlib maintainer). [^9] Using bors also enforces the not rocket science principle [16]: ensuring that the main (development) branch of Mathlib always passes all continuous integration tests. This policy has several reasons. For one, formalisation in Lean ensures a strong level of correctness: a formalised mathematical proof which compiles without any and is not intentionally manipulating the Lean environment or adding additional axioms—both of which are checked for by Mathlib’s continuous integration will be a correct and complete proof. However, designing an integrated mathematical library well still requires expert knowledge: usually, there are different potential designs, with vastly different usability and ergonomics. Code review is one way to form and share common knowledge about design best practices. It also has an empowering function: new contributors can propose new code, without being afraid of breaking other proofs (if their code compiles, the proofs are still fine) or overlooking important aspects—because an experienced Lean user will look over their contribution. This way, code review is a way of teaching and onboarding new contributors. Mathlib is growing steadily, as is its influx of new contributions. At the same time, many of its contributors are volunteers or work on Mathlib as a side project. Most of Mathlib’s maintainers have teaching, administrative or supervision commitments: allowing Mathlib to scale requires lightening their work. This has both social and technical aspects. On the social side, maintaining Mathlib means embracing teamwork and delegating responsibility. Approval of Mathlib pull requests is a multi-stage system, allowing many people to contribute. Code review in Mathlib is decentralised: everybody is allowed and encouraged to review contributions; there are guidelines teaching what to review for. This ranges from more basic aspects (such as adherence to Mathlib’s style guidelines, or whether new documentation is comprehensible) to advanced aspects like the correct code location and library design patterns. There are regular review meetings, where reviewers and interested contributors review code together—both to make this work more enjoyable, and to mentor potential new reviewers. Mathlib’s maintainers have the final say on approving a pull request. The number of maintainers keeps growing steadily, and is currently around 30. Maintainers have different specialisations; together, they cover the breadth of Mathlib. Having multiple knowledgeable maintainers in one area is beneficial, so they can discuss difficult decisions. Between maintainers and all contributors is the group of Mathlib reviewers. These are Mathlib contributors who have provided useful reviews in the past. A reviewer can approve a pull request by commenting maintainer merge, which notifies the maintainers. Usually, these get merged (or at least discussed further) more quickly. This system is a bit like journal referees and editors: the reviewer group lightens the load on maintainers; the direct notification mechanism provides an additional small incentive for reviewers. This works well in practice: in the last year, about one in three merged pull requests was approved by a reviewer before they were merged. On a technical level, we can save reviewer and maintainer time by automating tedious or mechanical aspects of this process. Two basic means of automation are using continuous integration with bors and automatic maintainer notifications through the maintainer merge command. Let us discuss Mathlib’s review automation from two further angles: tooling for reviewers and tooling for triage.
@@ 7.1 Tooling for reviewers
There are multiple ways for automation to support pull request reviewers. To begin with, we need to find a suitable reviewer for a given pull request. A pull request with a new theorem is best reviewed by somebody familiar with the theorem’s area; something similar holds for a contribution improving Mathlib’s infrastructure, adding a new linter or adding a new tactic. Each new pull request adding new content is automatically labelled according to its area. There is a file collecting each reviewer’s area of interest and expertise. This allows automatically suggesting suitable reviewers for every pull request—taking their expertise and reviewing capacity into account. At the time of writing, this automatic assignment is run on an experimental basis. This labelling information has also been used to identify missing reviewers: if pull requests in a certain area take a long time to get reviewed, recruiting further reviewers could be worthwhile. Secondly, we have created custom tooling exposing important information to reviewers. When a pull request is filed, a script creates a sticky comment summarising key information about a PR. This comment is updated in-place as the pull request is amended. The comment shows which lemmas were renamed (so reviewers can spot those more easily, and check if deprecation warnings are present). It also exposes when transitive imports of files change, which is extremely helpful for reviewing import refactoring. Finally, the summary comment shows any changes in technical debt metrics. Further, there is substantial automation to help with information flow within the project. Collaboration on Mathlib occurs primarily on two sites: GitHub, where git repositories including Mathlib and its dependencies are hosted, and the leanprover-community Zulip, a moderated chat site where community members discuss their contributions and other Lean-related topics. Automation can help with bridging these two venues and keeping some information synchronised. Whenever a specific pull request is mentioned in the Zulip chat, a script posts an emoji reaction indicating the current pull request status (such as, whether it is waiting for author changes, a review or was already merged). This allows reviewers and maintainers to see at a glance if action has already been taken. This reaction is updated by a GitHub actions script whenever the pull request’s status changes. We make extensive use of Zulip’s custom linkifier feature to allow users to link directly to issues or pull requests in the Mathlib repository, Mathlib documentation entries, and other useful pages.
@@ 7.2 Triage tooling
The number of open Mathlib pull requests is growing steadily. At the time of writing, there are about 1500 open pull requests. Often, new pull requests are getting reviewed quickly, but their sheer volume makes it difficult to keep track of every single PR. Better tooling helps to ensure that every pull request is kept track of and receives a timely response. Investigating this closely shows that different users of such tooling have differing needs. For instance, a reviewer may want to find reviewable pull requests matching their areas of interest. Mathlib’s maintainers would like to know about pull requests which were already approved by a reviewer. Other contributors may be interested in recent pull requests to “see what is going on”, or viewing all pull requests related to a particular area. Pull request authors may want to see the status of all their pull requests or find out why their pull request is not on the review queue. Most importantly, a triage team (see below) keeping track of pull requests’ state needs better tooling to see the state of each pull request. GitHub’s default interface can do many of the above tasks; however, at Mathlib’s scale we have increasingly run into its limitations. For example, to triage reviewable pull requests one would like to know the total time a pull request is already waiting for review. This need not be a continuous time span, so computing this needs to consider a pull request’s change over time. GitHub only exposes a “last updated” date, which does not provide this. (In addition, this “last update” measures almost every activity, which is often too sensitive for our purposes: a bystander’s comment or a notification of a merge conflict resets the last activity counter, but need not mean review activity.) Similarly, sorting pull requests by their total diff can be useful information, as is the number of PRs depending on a given PR. We have built a custom editorial tool to address all these shortcomings. Pull requests are grouped by their status (such as waiting for review, blocked on another pull requests or waiting for the author to respond to review comments). These categories are tailored to Mathlib’s infrastructure (for instance, pull request failures related to Mathlib’s infrastructure are highlighted automatically, speeding up the diagnosis and fix). In each category, we compactly display (see Figure 1) all pull requests together with additional meta-information such as its author, title and labels, but also additional information such as diff size or the total time in review. These lists can be semantically sorted by each category. This tooling can adapt and grow with Mathlib’s needs easily.
@@
We ran a survey among Mathlib contributors, reviewers and maintainers. The results provided many useful suggestions for improving this tool further (many of which have already been implemented). The survey result also affirms that this changes how contributors look for pull requests: the dashboard has become part of many maintainers’ standard workflow. The triage functionality in particular is used by maintainers, but also by the newly-founded triage team: having a common webpage to look at makes this process a lot smoother. Under the hood, this dashboard has two components. One keeps a cache, containing the metadata we care about for every open pull request. Periodically, we query GitHub’s graphql API to update this data for any pull requests which received updates, in effect creating a database of each pull request’s relevant state. The code for this amounts to a few hundred lines of shell and Python scripts: this could easily be run on a local computer, coupled with a cronjob. The other, user-facing component, processes this metadata to extract the aggregate data we care about. (For instance, from the history of all label changes on a pull request, we determine the total time that PR was ready for review.) Using this metadata, we create the triage dashboard. The dashboard is a static webpage, using jquery datatables to create the dashboards in each category: these can be sorted and filtered dynamically. Using a static webpage means fast response times (filtering a list of pull requests is almost instantaneous), but means the webpage needs to be manually refreshed to see updated changes. Altogether, this project consists of about 4500 lines of Python code, and about 1000 lines of shell scripts, .graphql schemas etc.
@@ 8 Conclusion
Growing Mathlib would be impossible without automation to scale its maintainance: this applies on the social, organisational and technical levels. While we cannot measure its effects precisely, the fact that Mathlib still grows fast shows how useful it is. For instance, Mathlib’s deprecation system has made updating to a newer version of Mathlib a much smoother process. Currently, such updates still require manual work, but this task can be automated well: both for deprecated declarations and modules, there are prototype tools performing the replacement automatically. An auto-formatter for Lean would improve dealing with code style questions significantly. Lean’s extensibility makes this difficult in general: users can define new syntax and custom notation, which can drastically change the look and feel of the language. [^10] For example, one can allow users to write in a form of controlled natural language [23]: formatting that would be a very different endavour. Formatting user-defined syntax and notation requires knowing the preferred formatting for these. While Mathlib makes extensive use of custom notation to make its code resemble mathematical notation, this is relatively benign: all syntax has an associated pretty-printer which indicates some reasonable formatting. Mathlib’s review and triage dashboard is mostly Mathlib-agnostic: the particular PR categories are tuned to Mathlib, but the underlying logic is not. With appropriate configuration, other open source projects could use this as well.
@@ Acknowledgements
We would like to thank the anonymous referees for their helpful comments and suggestions. Michael Rothgang was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813.

### v2

@@ 7 Scaling code review
Since its beginning, the Mathlib library required code review prior to merging changes. This is enforced through the bors bot: pull requests are merged automatically upon approval by an authorised user (i.e., a Mathlib maintainer). [^9] Using bors also enforces the not rocket science principle [16]: ensuring that the main (development) branch of Mathlib always passes all continuous integration tests.
This policy has several reasons. For one, formalisation in Lean ensures a strong level of correctness: a formalised mathematical proof which compiles without any and is not intentionally manipulating the Lean environment or adding additional axioms—both of which are checked for by Mathlib’s continuous integration will be a correct and complete proof. However, designing an integrated mathematical library well still requires expert knowledge: usually, there are different potential designs, with vastly different usability and ergonomics. Code review is one way to form and share common knowledge about design best practices. It also has an empowering function: new contributors can propose new code, without being afraid of breaking other proofs (if their code compiles, the proofs are still fine) or overlooking important aspects—because an experienced Lean user will look over their contribution. This way, code review is a way of teaching and onboarding new contributors.
Mathlib is growing steadily, as is its influx of new contributions. At the same time, many of its contributors are volunteers or work on Mathlib as a side project. Most of Mathlib’s maintainers have teaching, administrative or supervision commitments: allowing Mathlib to scale requires lightening their work. This has both social and technical aspects. On the social side, maintaining Mathlib means embracing teamwork and delegating responsibility. Approval of Mathlib pull requests is a multi-stage system, allowing many people to contribute.
Code review in Mathlib is decentralised: everybody is allowed and encouraged to review contributions; there are guidelines teaching what to review for. This ranges from more basic aspects (such as adherence to Mathlib’s style guidelines, or whether new documentation is comprehensible) to advanced aspects like the correct code location and library design patterns. There are regular review meetings, where reviewers and interested contributors review code together—both to make this work more enjoyable, and to mentor potential new reviewers.
Mathlib’s maintainers have the final say on approving a pull request. The number of maintainers keeps growing steadily, and is currently around 30. Maintainers have different specialisations; together, they cover the breadth of Mathlib. Having multiple knowledgeable maintainers in one area is beneficial, so they can discuss difficult decisions.
Between maintainers and all contributors is the group of Mathlib reviewers. These are Mathlib contributors who have provided useful reviews in the past. A reviewer can approve a pull request by commenting maintainer merge, which notifies the maintainers. Usually, these get merged (or at least discussed further) more quickly. This system is a bit like journal referees and editors: the reviewer group lightens the load on maintainers; the direct notification mechanism provides an additional small incentive for reviewers. This works well in practice: in the last year, about one in three merged pull requests was approved by a reviewer before they were merged.
On a technical level, we can save reviewer and maintainer time by automating tedious or mechanical aspects of this process. Two basic means of automation are using continuous integration with bors and automatic maintainer notifications through the maintainer merge command. Let us discuss Mathlib’s review automation from two further angles: tooling for reviewers and tooling for triage.
@@ 7.1 Tooling for reviewers
There are multiple ways for automation to support pull request reviewers. To begin with, we need to find a suitable reviewer for a given pull request. A pull request with a new theorem is best reviewed by somebody familiar with the theorem’s area; something similar holds for a contribution improving Mathlib’s infrastructure, adding a new linter or adding a new tactic. Each new pull request adding new content is automatically labelled according to its area. There is a file collecting each reviewer’s area of interest and expertise. This allows automatically suggesting suitable reviewers for every pull request—taking their expertise and reviewing capacity into account. At the time of writing, this automatic assignment is run on an experimental basis. This labelling information has also been used to identify missing reviewers: if pull requests in a certain area take a long time to get reviewed, recruiting further reviewers could be worthwhile.
Secondly, we have created custom tooling exposing important information to reviewers. When a pull request is filed, a script creates a sticky comment summarising key information about a PR. This comment is updated in-place as the pull request is amended. The comment shows which lemmas were renamed (so reviewers can spot those more easily, and check if deprecation warnings are present). It also exposes when transitive imports of files change, which is extremely helpful for reviewing import refactoring. Finally, the summary comment shows any changes in technical debt metrics.
Further, there is substantial automation to help with information flow within the project. Collaboration on Mathlib occurs primarily on two sites: GitHub, where git repositories including Mathlib and its dependencies are hosted, and the leanprover-community Zulip, a moderated chat site where community members discuss their contributions and other Lean-related topics. Automation can help with bridging these two venues and keeping some information synchronised. Whenever a specific pull request is mentioned in the Zulip chat, a script posts an emoji reaction indicating the current pull request status (such as, whether it is waiting for author changes, a review or was already merged). This allows reviewers and maintainers to see at a glance if action has already been taken. This reaction is updated by a GitHub actions script whenever the pull request’s status changes.
We make extensive use of Zulip’s custom linkifier feature to allow users to link directly to issues or pull requests in the Mathlib repository, Mathlib documentation entries, and other useful pages.
@@ 7.2 Triage tooling
The number of open Mathlib pull requests is growing steadily. At the time of writing, there are about 1500 open pull requests. Often, new pull requests are getting reviewed quickly, but their sheer volume makes it difficult to keep track of every single PR. Better tooling helps to ensure that every pull request is kept track of and receives a timely response.
Investigating this closely shows that different users of such tooling have differing needs. For instance, a reviewer may want to find reviewable pull requests matching their areas of interest. Mathlib’s maintainers would like to know about pull requests which were already approved by a reviewer. Other contributors may be interested in recent pull requests to “see what is going on”, or viewing all pull requests related to a particular area. Pull request authors may want to see the status of all their pull requests or find out why their pull request is not on the review queue. Most importantly, a triage team (see below) keeping track of pull requests’ state needs better tooling to see the state of each pull request.
GitHub’s default interface can do many of the above tasks; however, at Mathlib’s scale we have increasingly run into its limitations. For example, to triage reviewable pull requests one would like to know the total time a pull request is already waiting for review. This need not be a continuous time span, so computing this needs to consider a pull request’s change over time. GitHub only exposes a “last updated” date, which does not provide this. (In addition, this “last update” measures almost every activity, which is often too sensitive for our purposes: a bystander’s comment or a notification of a merge conflict resets the last activity counter, but need not mean review activity.) Similarly, sorting pull requests by their total diff can be useful information, as is the number of PRs depending on a given PR.
We have built a custom editorial tool to address all these shortcomings. Pull requests are grouped by their status (such as waiting for review, blocked on another pull request or waiting for the author to respond to review comments). These categories are tailored to Mathlib’s infrastructure (for instance, pull request failures related to Mathlib’s infrastructure are highlighted automatically, speeding up the diagnosis and fix). In each category, we compactly display (see Figure 1) all pull requests together with additional meta-information such as its author, title and labels, but also additional information such as diff size or the total time in review. These lists can be semantically sorted by each category. This tooling can adapt and grow with Mathlib’s needs easily.
@@
We ran a survey among Mathlib contributors, reviewers and maintainers. The results provided many useful suggestions for improving this tool further (many of which have already been implemented). The survey result also affirms that this changes how contributors look for pull requests: the dashboard has become part of many maintainers’ standard workflow. The triage functionality in particular is used by maintainers, but also by the newly-founded triage team: having a common webpage to look at makes this process a lot smoother.
Under the hood, this dashboard has two components. One keeps a cache, containing the metadata we care about for every open pull request. Periodically, we query GitHub’s graphql API to update this data for any pull requests which received updates, in effect creating a database of each pull request’s relevant state. The code for this amounts to a few hundred lines of shell and Python scripts: this could easily be run on a local computer, coupled with a cronjob.
The other component is responsible for the final user interface. It processes the downloaded metadata to extract the aggregate data we care about. (For instance, from the history of all label changes on a pull request, we determine the total time that PR was ready for review.) Using this metadata, we create the triage dashboard. The dashboard is a static webpage, using jQuery datatables to create the dashboards in each category: these can be sorted and filtered dynamically. Using a static webpage means fast response times (filtering a list of pull requests is almost instantaneous), but means the webpage needs to be manually refreshed to see updated changes. Altogether, this project consists of about 4500 lines of Python code, and about 1000 lines of shell scripts, .graphql schemas etc.
@@ 8 Conclusion
Growing Mathlib would be impossible without automation to scale its maintainance: this applies on the social, organisational and technical levels. While we cannot measure its effects precisely, the fact that Mathlib still grows fast shows how useful it is. For instance, Mathlib’s deprecation system has made updating to a newer version of Mathlib a much smoother process. Currently, such updates still require manual work, but this task can be automated well: both for deprecated declarations and modules, there are prototype tools performing the replacement automatically.
An auto-formatter for Lean would improve dealing with code style questions significantly. Lean’s extensibility makes this difficult in general: users can define new syntax and custom notation, which can drastically change the look and feel of the language. [^10] For example, one can allow users to write in a form of controlled natural language [24]: formatting that would be a very different endavour. Formatting user-defined syntax and notation requires knowing the preferred formatting for these. While Mathlib makes extensive use of custom notation to make its code resemble mathematical notation, this is relatively benign: all syntax has an associated pretty-printer which indicates some reasonable formatting.
Mathlib’s review and triage dashboard is mostly Mathlib-agnostic: the particular PR categories are tuned to Mathlib, but the underlying logic is not. With appropriate configuration, other open source projects could use this as well.
@@ Acknowledgements
We would like to thank the anonymous referees for their helpful comments and suggestions. Michael Rothgang was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813. This material is based upon work supported by the U. S. National Science Foundation Award No. DMS-2302263.